# Equilibrium Policy Generalization: A Reinforcement Learning Framework for Cross-Graph Zero-Shot Generalization in Pursuit-Evasion Games

### v1

@@ 4 Experiments
In this section, we verify that our EPG framework can train a generalized policy with robust zero-shot performance under unseen graph structures. Using $76$ procedurally generated maps in the Dungeon environment [4], we construct a heterogeneous training set by discretizing each map into small-scale (100-node) and large-scale (500-node) graphs. The policy training follows the cross-graph reinforcement learning pipeline in Section 3.1. In no-exit scenarios with $2$ or $3$ pursuers, we directly use the DP algorithm (Algorithm 1) in Section 3.2 as an accurate equilibrium oracle. When there are more pursuers, we use the grouping extension of the DP approach to construct an approximate equilibrium oracle. In multi-exit scenarios, we use the heuristic approach to construct the approximate oracle. Besides, we use the sequence policy model in Section 3.3 to represent either pursuer or evader policy. For the evader side, the policy model is reduced to a single policy network, with $\mu^{*}$ and $\nu^{*}$ exchanged in Figure 1. The training time costs are reported in Appendix C.3. We test the zero-shot performance of the learned pursuer or evader policy under unseen graphs without further fine-tuning.
@@ 4.3 Performance Comparisons in Graphs with Exits
Figure 4: Zero-shot performance in 55-exit PEGs during training (left) and testing (mid & right)
Since the state-of-the-art approaches in graph-based PEGs (including Grasper [16] and multi-task PSRO (MT-PSRO) [17]) deal with multi-exit PEGs, we also apply our EPG framework to the same problem setting of $5$ pursuers, $1$ evader, and $8$ exits described in the paper of Grasper [16]. As is mentioned in Section 3.2, we design an equilibrium heuristic based on bipartite graph matching (see Appendix B) to construct an approximate equilibrium oracle. As is shown in Figure 4 (left), our EPG framework steadily trains the RL pursuers under different hyperparameters $\beta\in\{0.01,0.1\}$ when using the heuristic approach. Mere supervised learning (green) without SAC loss can no longer guarantee a steady improvement in unseen graphs, though it can simply work in the no-exit scenario.
Since [16] also uses Grid Map and Scotland-Yard Map as two fixed testing graphs, it is direct to evaluate the zero-shot performance of our RL pursuers under the same game rules. The comparative methods of Grasper, MT-PSRO, and MT-PSRO with augmentation (MT-PSRO-Aug) all train their policies given the graph structure. Nevertheless, Figure 4 (mid & right; with results from [16]) shows that our zero-shot performance (the dashed lines) under the unseen graphs ($0.637$ for Grid Map and $0.652$ for Scotland-Yard Map) significantly outperforms their zero-shot performance (the starting points), which is with respect to unseen initial conditions. Even after a $10$-minute fine-tuning process using PSRO [15], only Grasper in the Scotland-Yard Map can match the zero-shot performance of our policy. This result further demonstrates the generalization capability of our EPG framework.
@@ Appendix B Equilibrium Heuristic for PEGs with Exits
In no-exit graphs, the termination function $f(s_{p},s_{e})$ is $1$ when a number of pursuers are adjacent to the evader. Since there is no termination condition about exits, the pursuers only care about how soon they can capture the evader. In graphs with exits, however, the pursuers lose when $g(s_{p},s_{e})=1$, and a successful pursuit requires strict overlapping rather than adjacency (see [16]). Therefore, to block the exits around the evader is more reasonable than to block all possible actions of the evader and capture it. For the evader, it is in turn reasonable to select an exit that cannot be blocked in time.
@@ C.1 Implementation Details
For the training of Q functions, which is not specified in Section 3.3, we employ the common technique of double target networks to avoid overestimations [12]. When there are two or three pursuers, we simply use centralized value networks to directly represent Q functions. In the scenarios with more pursuers, we use value-decomposition networks [26] as a simple way to decompose the joint Q functions. As the pursuers are homogeneous agents, we have a direct decomposition ${{Q}_{\phi}}(s,a)=\sum\nolimits_{i=1}^{m}{{{Q}_{\phi}}(s,{{a}_{i}})}$, where $m$ is the number of the pursuers.
In either the no-exit or multi-exit scenario, we use the same hyperparameter setting shown in Table 3 throughout the EPG training for either pursuer or evader policy. Note that the target entropy $\overline{H}$ in SAC [5] is defined in the form of $\log\left(1/\left|\mathcal{A}\right|\right)=\log\left|\mathcal{A}\right|$ multiplied by a predetermined coefficient.
@@ C.3 Compute Resources and Time Costs
5-pursuer 8-exit PEG [16] 5-pursuer policy $159$ min
@@ D.2 Test Details
For multi-exit PEGs, we randomly generated $1000$ initial positions for the exits and agents, strictly following the descriptions and code from Grasper [16] for Grid Map and Scotland-Yard Map. For example, the minimum length of the evader’s shortest path to any exit is set to 6 for the Grid Map and 5 for the Scotland-Yard Map. While [16] claims that they rule out the trivial cases that are either too difficult or too simple, the specific restrictions on the initial conditions are related to their method and hard to replicate. Therefore, we simply rule out the cases where it is clearly impossible for the pursuers or the evader to win. For the evader, we require that its distance to the closest exit be no greater than $10$ (the time horizon of the game). For the pursuers, we require that, for any exit that the evader can reach in $10$ steps, there is at least one pursuer who is closer than or as close as the evader with respect to the shortest path distance towards the exit.
@@ D.4 Pursuit Examples in Multi-Exit Graphs
While the time horizon for no-exit PEGs is $128$, it is only set to $10$ in multi-exit PEGs to align with [16]. Therefore, we show more pursuit examples in graphs with exits from the Dungeon environment. Figures 13, 14, and 15 provide three examples to show the zero-shot performance of our $5$-pursuer policy trained through EPG with equilibrium heuristic. The light stars represent the $8$ exits.
@@ E.1 Results in Scale-Free Graph
Figure 16 provides an illustration of Scale-Free Graph, which is generated using Barabási-Albert preferential attachment [1]. With a maximum degree of $21$ and a diameter of $11$, Scale-Free Graph has very strong connectivity that can make evasion easier. As the pursuit is not guaranteed for $2$ pursuers in this graph, we find that the DP pursuer policy only has a success rate of $0.1$ against the DP evader. With a success rate of $0.44$, our RL policy clearly outperforms DP under this graph.
@@ F.2 Performance against PSRO Policies
Besides DP, we also test the results of our EPG pursuer and evader policies against the PSRO [15] policies directly trained on the testing graphs. As is shown in Table 8, our EPG policies guarantee to capture the PSRO evader and survive significantly longer against the PSRO pursuers (except for Grid Map). This further verifies the benefit of using EPG for cross-graph zero-shot generalization.

### v2

@@ 4 Experiments
In this section, we verify that our EPG framework can train a generalized policy with robust zero-shot performance under unseen graph structures. Using $76$ procedurally generated maps in the Dungeon environment [chen2019self], we construct a heterogeneous training set by discretizing each map into small-scale (100-node) and large-scale (500-node) graphs. The policy training follows the cross-graph reinforcement learning pipeline in Section 3.1. In no-exit scenarios with $2$ or $3$ pursuers, we directly use the DP algorithm (Algorithm 1) in Section 3.2 as an accurate equilibrium oracle. When there are more pursuers, we use the grouping extension of the DP approach to construct an approximate equilibrium oracle. In multi-exit scenarios, we use the heuristic approach to construct the approximate oracle. Besides, we use the sequence policy model in Section 3.3 to represent either pursuer or evader policy. For the evader side, the policy model is reduced to a single policy network, with $\mu^{*}$ and $\nu^{*}$ exchanged in Figure 1. The training time costs are reported in Appendix C.3. We test the zero-shot performance of the learned pursuer or evader policy under unseen graphs without further fine-tuning.
@@ 4.3 Performance Comparisons in Graphs with Exits
Figure 4: Zero-shot performance in 88-exit PEGs during training (left) and testing (mid & right)
Since the state-of-the-art approaches in graph-based PEGs (including Grasper [li2024grasper] and multi-task PSRO (MT-PSRO) [li2023solving]) deal with multi-exit PEGs, we also apply our EPG framework to the same problem setting of $5$ pursuers, $1$ evader, and $8$ exits described in the paper of Grasper [li2024grasper]. As is mentioned in Section 3.2, we design an equilibrium heuristic based on bipartite graph matching (see Appendix B) to construct an approximate equilibrium oracle. As is shown in Figure 4 (left), our EPG framework steadily trains the RL pursuers under different hyperparameters $\beta\in\{0.01,0.1\}$ when using the heuristic approach. Mere supervised learning (green) without SAC loss can no longer guarantee a steady improvement in unseen graphs, though it can simply work in the no-exit scenario.
Since [li2024grasper] also uses Grid Map and Scotland-Yard Map as two fixed testing graphs, it is direct to evaluate the zero-shot performance of our RL pursuers under the same game rules. The comparative methods of Grasper, MT-PSRO, and MT-PSRO with augmentation (MT-PSRO-Aug) all train their policies given the graph structure. Nevertheless, Figure 4 (mid & right; with results from [li2024grasper]) shows that our zero-shot performance (the dashed lines) under the unseen graphs ($0.637$ for Grid Map and $0.652$ for Scotland-Yard Map) significantly outperforms their zero-shot performance (the starting points), which is with respect to unseen initial conditions. Even after a $10$-minute fine-tuning process using PSRO [lanctot2017unified], only Grasper in the Scotland-Yard Map can match the zero-shot performance of our policy. This result further demonstrates the generalization capability of our EPG framework.
@@ Appendix B Equilibrium Heuristic for PEGs with Exits
In no-exit graphs, the termination function $f(s_{p},s_{e})$ is $1$ when a number of pursuers are adjacent to the evader. Since there is no termination condition about exits, the pursuers only care about how soon they can capture the evader. In graphs with exits, however, the pursuers lose when $g(s_{p},s_{e})=1$, and a successful pursuit requires strict overlapping rather than adjacency (see [li2024grasper]). Therefore, to block the exits around the evader is more reasonable than to block all possible actions of the evader and capture it. For the evader, it is in turn reasonable to select an exit that cannot be blocked in time.
@@ C.1 Implementation Details
For the training of Q functions, which is not specified in Section 3.3, we employ the common technique of double target networks to avoid overestimations [hasselt2010double]. When there are two or three pursuers, we simply use centralized value networks to directly represent Q functions. In the scenarios with more pursuers, we use value-decomposition networks [sunehag2017value] as a simple way to decompose the joint Q functions. As the pursuers are homogeneous agents, we have a direct decomposition ${{Q}_{\phi}}(s,a)=\sum\nolimits_{i=1}^{m}{{{Q}_{\phi}}(s,{{a}_{i}})}$, where $m$ is the number of the pursuers.
In either the no-exit or multi-exit scenario, we use the same hyperparameter setting shown in Table 3 throughout the EPG training for either pursuer or evader policy. Note that the target entropy $\overline{H}$ in SAC [christodoulou2019soft] is defined in the form of $\log\left(1/\left|\mathcal{A}\right|\right)=\log\left|\mathcal{A}\right|$ multiplied by a predetermined coefficient.
@@ C.3 Compute Resources and Time Costs
5-pursuer 8-exit PEG [li2024grasper] 5-pursuer policy $159$ min
@@ D.2 Test Details
For multi-exit PEGs, we randomly generated $1000$ initial positions for the exits and agents, strictly following the descriptions and code from Grasper [li2024grasper] for Grid Map and Scotland-Yard Map. For example, the minimum length of the evader’s shortest path to any exit is set to 6 for the Grid Map and 5 for the Scotland-Yard Map. While [li2024grasper] claims that they rule out the trivial cases that are either too difficult or too simple, the specific restrictions on the initial conditions are related to their method and hard to replicate. Therefore, we simply rule out the cases where it is clearly impossible for the pursuers or the evader to win. For the evader, we require that its distance to the closest exit be no greater than $10$ (the time horizon of the game). For the pursuers, we require that, for any exit that the evader can reach in $10$ steps, there is at least one pursuer who is closer than or as close as the evader with respect to the shortest path distance towards the exit.
@@ D.4 Pursuit Examples in Multi-Exit Graphs
While the time horizon for no-exit PEGs is $128$, it is only set to $10$ in multi-exit PEGs to align with [li2024grasper]. Therefore, we show more pursuit examples in graphs with exits from the Dungeon environment. Figures 13, 14, and 15 provide three examples to show the zero-shot performance of our $5$-pursuer policy trained through EPG with equilibrium heuristic. The light stars represent the $8$ exits.
@@ E.1 Results in Scale-Free Graph
Figure 16 provides an illustration of Scale-Free Graph, which is generated using Barabási-Albert preferential attachment [barabasi1999emergence]. With a maximum degree of $21$ and a diameter of $11$, Scale-Free Graph has very strong connectivity that can make evasion easier. As the pursuit is not guaranteed for $2$ pursuers in this graph, we find that the DP pursuer policy only has a success rate of $0.1$ against the DP evader. With a success rate of $0.44$, our RL policy clearly outperforms DP under this graph.
@@ F.2 Performance against PSRO Policies
Besides DP, we also test the results of our EPG pursuer and evader policies against the PSRO [lanctot2017unified] policies directly trained on the testing graphs. As is shown in Table 8, our EPG policies guarantee to capture the PSRO evader and survive significantly longer against the PSRO pursuers (except for Grid Map). This further verifies the benefit of using EPG for cross-graph zero-shot generalization.